Problem G.11 Maria Miller is a sales executive at a Baltimore tirm. She is 25 years old and plans to invest $4,400 each year in an IRA account until she is 65 at which time she will retire (a total of 10 payments). If Maria invests at the beginning of each year, and the IRA investment will earn 10.10 percent annually, how much will she have when she retires? Assume that she makes the first payment today. (Round factor values to 4 decimal places, e.g. 1.5212 and fial answer to 2 decimal places, e.g. 15.25.) Future value of investment
Answer #1

P = annual payments = 4400

n = number of payments = 40

interest rate = r = 10.10% = 0.1010

We need to calculate the amount accumulated after 40 years, i.e. Future Value (FV) of the investment

Hence, FV = P(1+r)n + P(1+r)n-1 +....+ P(1+r) = P [((1 + r)n - 1) / r])(1 + r) = 4400 [(1.101040 - 1)/0.1010](1.1010)

=> FV = $2,203,221.65